OpenStudy (anonymous):

Solve logx + 5 36 = 2 Please help me understand what "log" means and what it does to the equation?

OpenStudy (anonymous):

\[\log_{x+5}(36)=2\] means in equivalent exponential form that \[(x+5)^2=36\] now solve for x

OpenStudy (anonymous):

you change \[\log_b(x)=y\] to \[b^y=x\] as equivalent exponential form

OpenStudy (anonymous):

now it is easy enough to solve. write \[x+5=6\] or \[x+5=-6\] so \[x=1\] or \[x=-11\] but the base cannot be negative so it is \[x=6\]

OpenStudy (anonymous):

sorry i mean \[x=1\] not 6

OpenStudy (anonymous):

\[x=1\] so \[x+5=6\] is the base

OpenStudy (anonymous):

But if it could be negative would it be -6?

OpenStudy (anonymous):

or 6?

OpenStudy (anonymous):

Wait can it ever be negative?